Calculate the pH for 100. mL of a 0.500 M solution of ammonia, (NH3   Kb=1.8x10-5) being titrated...

  1. Calculate the pH for 100. mL of a 0.500 M solution of ammonia, (NH3   Kb=1.8x10-5) being titrated with 0.500 M HCl at the following positions in the titration.

    i) The initial pH (before any HCl has been added).

    A.

    9.44

    B.

    10.81

    C.

    11.48

    D.

    12.00

    E.

    11.75
Answer #1

Ammonia is weak base dissociate as

NH3 + H2O   phpWoaLWa.png NH4+ + OH-

Kb = [NH4+] [OH-] / [NH3]

but [NH4+] = [OH-]

consider [NH4+] = [OH-] = X

then Kb = [X] [X] / [NH3]

Kb = X2 / [NH3]

X2 = Kb X [NH3]

substitute the value

X2 = (1.8 X 10-5) X (0.500) = 9 X 10-6

X = 0.003 M

[NH4+] = [OH-] = X = 0.003 M

[OH-] = 0.003 M

pOH = -log [OH-] = -log 0.003 = 2.52

pH = 14 - pH = 14 - 2.52 = 11.48

pH of 0.500 M NH3 = 11.48

Ans = C) 11.48